whats the next 3 numbers in the sequence..
4, 7, 15, 28, 46 ...

Answer by texttutoring(324)

Every time, the amount the next number goes up by is 5 more than the one before it.
For instance, 4 to 7 jumped up by 3. Then 7 to 15 jumped up by 8 (3+5=8)
15 to 28 jumped by 13 (8+5=13), and 28 to 46 jumped by 18 (13+5=18)

So for the next one:

46 will jump up to (18+5=23) 69
69 will jump up to (23+5=28) 97
97 will jump up to (28+5=33) 130

So the next 3 numbers are: 69,97,130.

Hope that makes sense!